Given that log4=0.6020 and log3=0.771, use the laws of logathrimns to find the log0.75 and log81

Found 4 solutions by Alan3354, Edwin McCravy, ikleyn, greenestamps:
Answer by Alan3354(69443)

Given that log4=0.6020 and log3=0.771, use the laws of logathrimns to find the log0.75 and log81
---------------------
log(3) is not 0.771

Answer by Edwin McCravy(20055)

log(0.75) = log(75/100) = log(3/4) = log(3)-log(4) = you finish

log(81) = log(34) = 4log(3) = you finish

Edwin

Answer by ikleyn(52781)

.
Given that log4=0.6020 and log3=0.771, use the laws of logathrimns to find the log0.75 and log81
~~~~~~~~~~~~~~~~~~~~~~~~


The problem's formulation in your post is  FATALLY  WRONG.

log(3)  is definitely less than  log(4),  but your post,  by mistake,
provides the value of  log(3)  GREATER  than  log(4).



Answer by greenestamps(13200)



Pay attention to what you are posting before you post it.

log(3) is 0.4771, not 0.771.

Then, using basic rules for logarithms....

log(0.75) = log(3/4) = log(3) - log(4)

log(81) = log(3^4) = 4*log(3)

You can do the arithmetic to get the answers.